# Multiverse Through Deepfakes: The MultiFakeVerse Dataset of Person-Centric Visual and Conceptual Manipulations

### v1

@@ 2. Related Work
Deepfake Datasets. The current landscape of image-based deepfake datasets is shown in Table 1. Most existing deepfake datasets focus primarily on person-level changes, where manipulations are confined to a person’s face or facial expressions. A prominent example is DFFD (Stehouwer et al., 2019), which contains 300K samples, which are generated and/or edited facial images with GAN methods. Other prominent datasets, which have been commonly used are FFHQ (Karras et al., 2019) and FakeSpotter  (Wang et al., 2019). Moving beyond faces, OHImg (May et al., 2023) introduced a deepfake dataset focused on overhead aerial imagery, while M3Dsynth (Zingarini et al., 2024) addressed medical image based manipulations. More recently, Huang et al. proposed the SIDA dataset (Huang et al., 2024), which leverages vision-language models to identify objects in images and captions, replacing them via inpainting with Stable Diffusion (e.g. cat $\rightarrow$ dog). However, these datasets primarily focus on object- and scene-level manipulations and do not address person-level edits that alter the perceived meaning of an image.
@@
Deepfake Detection. Deepfake detection techniques are commonly framed as classification problems within a data-centric framework (Sharma et al., 2024). These approaches often rely on a range of neural network architectures, notably Convolutional Neural Networks (CNNs) (Nguyen et al., 2019) and Transformers (Wang et al., 2024), to identify unique artifacts introduced during image based manipulation. A number of deepfake detection methods (Miao et al., 2023; Huang et al., 2024) have expanded beyond simple binary classification by developing datasets with pixel-level annotations of tampered regions, thereby supporting both detection and localization tasks. However, such resources are predominantly centered on facial forgeries (Miao et al., 2023), with limited availability of datasets targeting non-facial manipulations, and even fewer large-scale, publicly accessible datasets that reflect the diversity of scene content. To bridge these gaps, our work introduces a comprehensive dataset featuring a wide range of manipulations beyond facial edits, with a particular emphasis on realistic image-based manipulation and localization.
@@
LLM Powered Reasoning for Generation. The emergence of diffusion models has significantly advanced the field of image editing/manipulations (Wang et al., 2025). Recent progress in image inpainting, under both text-guided and unguided settings, has enabled more precise and high-quality manipulations (Wang et al., 2025). Text-conditioned inpainting (Xie et al., 2023) allows for detailed control over edits using natural language, while unconditioned methods offer strong baseline capabilities without explicit guidance. Traditionally, inpainting techniques (Huang et al., 2024) rely on the use of masks to specify regions for modification. In contrast, prompt-based image editing (Touvron et al., 2023a) allows for targeted alterations driven solely by text inputs, eliminating the need for explicit masks. Frameworks such as LANCE (Prabhu et al., 2023) and InstructPix2Pix (Brooks et al., 2022) harness this approach to automate image perturbation pipelines. LANCE (Prabhu et al., 2023), in particular, combines large language models (LLMs) with image captioning techniques to facilitate diverse, human-free image editing workflows. Expanding on this line of work, our method uses GPT-Image-1  (OpenAI, 2024), Gemini-2.0-Flash-Image-Generation (Anil and et al., 2023) and ICEdit (Zhang et al., 2025) to support image manipulation with a wider spectrum of perturbation intensities, guided by definitions of semantic change. To achieve this, we integrate models like Gemini-2.0-Flash (Anil and et al., 2023), ChatGPT-4o-latest (OpenAI, 2024) using prompt-based editing to generate precise, semantically meaningful image modifications. Multimodal models like Gemini, ChatGPT-4o, and Icedit leverage both text and image inputs to enable precise, context-aware edits and have a natural advantage over text-driven models like LANCE and InstructPix2Pix in fine-grained and semantically consistent image manipulation tasks. While text-driven models are effective for stylistic transformations, multimodal approaches offer broader capabilities and greater control over localized edits.
@@ 3.1.2. VLM based Image perception Manipulation
In this step, for each of the $<$referring expression, edit instruction$>$ pair, we prompt a VLM to perform the edit on the image, while ensuring that nothing else changes. We experiment with three different VLMs here - GPT-Image-1, Gemini-2.0-Flash-Image-Generation and ICEdit (Zhang et al., 2025). After observing  22K generated images, Gemini
@@
(1) the person you identified appears naive
(2) the person you identified appears nonchalant
(3) the person you identified appears proud
(4) the person you identified appears remorseful
(5) the person you identified appears inexperienced
(6) some factual information depicted in the image changes
Do not include any other information in the response.
emerged as the VLM for highest quality manipulations, as its edits are the most coherent with the rest of the scene with no changes to the untargeted regions of the input image.The ICEdit model often results in easily identifiable fakes with significant artifacts, whereas GPT-image-1 tends to edit in a few cases, untargeted regions due to constraints in the output aspect ratios (current options are: $1024\times 1024$, $1536\times 1024$, $1024\times 1536$).
3.2. Analyzing the Manipulated images
Ratio of edited area To obtain a measure of the spatial degree of modification, we calculate the mean squared difference between the pixel values of the original and altered images, threshold the difference to remove noise, and perform connected component analysis to obtain the masks of the edited regions. The ratio of edited area to the total image area is then plotted over the entire dataset, shown in Figure 3(a), which displays a distribution spread out from 0 to 0.8, thus signifying the wide variety of edits present in our dataset.
Image Captioning We use ShareGPT4V  (Chen et al., 2024) VLM model to obtain rich captions of both the original as well as tampered images. These
@@ 3.2. Analyzing the Manipulated images
are then used to calculate the directional similarity (Gal et al., 2022) between the real versus edited images and their corresponding captions as follows: ($CLIP_{OriginalImage}$ $CLIP_{TamperedImage}$) $\cdot$ ($CLIP_{OriginalImageCaption}$ $CLIP_{TamperedImageCaption}$). Particularly, we use Long-CLIP (Zhang et al., 2024) to calculate the image and text embeddings. From this analysis, we find that the maximum semantic change is observed in the object category, where the modification is around or on the person.
Level of manipulation performed We categorize the manipulations into Person-level, Object-level and Scene-level depending upon whether the manipulation target is a person (facial expression, identity, gaze, pose, or clothing), an object connected to a person (i.e. an object in the foreground - appearance change, addition, removal etc.) or some component in the overall scene/background of the image respectively. Since an image can have multiple levels of changes depending upon the edit instruction, we display the distribution of the these alteration levels in our dataset through a Venn diagram in Figure 2(b). The distribution shows that the different levels of manipulations are widely distributed in our dataset with around one-third of the edits being person-only, around one-fifth being scene-only and one-sixth being object-only.
Perceptual Impact of the edit on the viewer To obtain a measure of the change in the viewer perception as a result of the image tamperings, we consider the impact on the following six perceptual factors: (A) the viewer’s emotion or interpretation (B) the perceived identity or character of the person in the image (C) the perceived power dynamics/relationships (D) the perceived scene narrative (E) the possible intent of manipulation, and (F) the ethical implications of the tamperings. We prompt the Gemini-2.0-Flash VLM to describe these impacts, whose instruction details are shown in Prompt 3.2. The word clouds of the changes based on responses obtained from the VLM are shown in Figure 3. The first word map, corresponding to the changes in emotion perception (Figure 3(b)) features words such as significant, joyful, engaging, approachable, facial suggesting that our semantic modifications subtly influence the emotional perception. In case of scene context and narrative shifts (Fig. 3(c)), the word cloud contains terms such as focused, professional, potential, different, indicating plausible alterations to the conveyed story or setting due to the corresponding edits. Fig. 3(d), which illustrates the word map for shift in the perceived identity of depicted persons includes terms such asplayful, younger, vulnerable implying that our manipulations are able to influence the individual’s identity. The word map representing the intent behind the manipulations in Fig. 3(e) shows deceptive, persuasive and aesthetic as the most prominent words, reflecting a tendency for the manipulations to be viewed as intentional acts of persuasion or deception. According to the word map in Fig. 3(f), though the power dynamics and relationships remain stable for many cases, the presence of words like altered, dynamic, formal, competitive highlights that such dynamics do shift noticeably in a subset of the images. In Figure 4, we show some sample responses obtained for two such edits, with the different perceptual impacts highlighted. The VLM responses show that around 81% of the edits have mild ethical implications, while 14.2% have moderate and 0.3% of the edits have severe ethical implications in our dataset.
@@
Quality of generations. To evaluate the visual quality of MultiFakeVerse, we used the peak signal-to-noise ratio (PSNR), structural similarity index (SSIM) (Wang et al., 2004) and Fréchet inception distance (FID) (Heusel et al., 2017) metrics as shown in Table 3. SSIM value 0 indicates no similarity between the compared images, while 1 indicates perfect similarity, therefore, our value of $0.5774$ is indicative of the targeted edits which ensure that the untargeted regions remain as close as possible to the original image. Similarly, a low FID value of 3.30 indicates that the quality and diversity of the generated fakes is excellent. A high PSNR value of 66.30 dB is indicative of good image quality.
@@ 3.2.2. Computational Cost
4. Benchmarks and Experiments
Data Partitioning and Evaluation Metrics: We split the dataset into train, validation, and test sets as follows: first we randomly select 70% of the real images as train set, 10% real images as validation set and remaining 20% as test set. Next we add the images generated corresponding to each of the real images in the same set as the real image. We evaluate detection using image-level accuracy and F1 scores. For forgery localization, we use Area Under the Curve (AUC), F1 scores and Intersection over Union (IoU).
@@ 4. Benchmarks and Experiments
Detection Evaluation. We compare MultiFakeVerse against several SOTA deepfake detection methods on the test set, including CnnSpot (Wang et al., 2020), AntifakePrompt (Chang et al., 2024), TruFor (Guillaro et al., 2023) and VLM based SIDA (Huang et al., 2024). To ensure a fair comparison, we first evaluate these models on our dataset using their original pre-trained weights (zero-shot setting), and then retrain two of them (CNNSpot and SIDA) with our train dataset to assess performance improvements. Table 4 demonstrates that these models trained on earlier inpainting-based fakes struggle to identify our VLM-Editing based forgeries, particularly, CNNSpot tends to classify almost all the images as real. AntiFakePrompt has the best zero-shot performance with 66.87% average class-wise accuracy and 55.55% F1 score. After finetuning on our train set, we observe a performance improvement in both CNNSpot and SIDA-13B, with CNNSpot surpassing SIDA-13B in terms of both average class-wise accuracy (by 1.92%) as well as F1-Score (by 1.97%).
Localization Evaluation. We evaluate the localization performance of SIDA-13B (Huang et al., 2024) model on our test set, using the perturbation masks obtained in Section 3.2, both using their pre-trained weights as well as after finetuning on our dataset. In zero-shot setting, the model has an IoU of 13.10, F1-Score 19.92 and AUC 14.06. After finetuning on our train dataset, the localization performance of the SIDA-13B model remains at an IoU of 24.74, F1-Score 39.40 and AUC 37.53, signifying that the model struggles to accurately identify the manipulated image regions. Thus, there is room for further improvement in the forgery localization methods, with MultiFakeVerse as an ideal benchmark.

### v2

@@ 2. Related Work
2.0.1. Deepfake Datasets
The current landscape of image-based deepfake datasets is shown in Table 1. Most existing deepfake datasets focus primarily on person-level changes, where manipulations are confined to a person’s face or facial expressions. A prominent example is DFFD (Stehouwer et al., 2019), which contains 300K samples, which are generated and/or edited facial images with GAN methods. Other prominent datasets, which have been commonly used are FFHQ (Karras et al., 2019) and FakeSpotter  (Wang et al., 2019). Moving beyond faces, OHImg (May et al., 2023) introduced a deepfake dataset focused on overhead aerial imagery, while M3Dsynth (Zingarini et al., 2024) addressed medical image based manipulations. More recently, Huang et al. proposed the SIDA dataset (Huang et al., 2024), which leverages vision-language models to identify objects in images and captions, replacing them via inpainting with Stable Diffusion (e.g. cat $\rightarrow$ dog). However, these datasets primarily focus on object- and scene-level manipulations and do not address person-level edits that alter the perceived meaning of an image.
@@ 2.0.2. Deepfake Detection
Deepfake detection techniques are commonly framed as classification problems within a data-centric framework (Sharma et al., 2024). These approaches often rely on a range of neural network architectures, notably Convolutional Neural Networks (CNNs) (Nguyen et al., 2019) and Transformers (Wang et al., 2024), to identify unique artifacts introduced during image based manipulation. A number of deepfake detection methods (Miao et al., 2023; Huang et al., 2024) have expanded beyond simple binary classification by developing datasets with pixel-level annotations of tampered regions, thereby supporting both detection and localization tasks. However, such resources are predominantly centered on facial forgeries (Miao et al., 2023), with limited availability of datasets targeting non-facial manipulations, and even fewer large-scale, publicly accessible datasets that reflect the diversity of scene content. To bridge these gaps, our work introduces a comprehensive dataset featuring a wide range of manipulations beyond facial edits, with a particular emphasis on realistic image-based manipulation and localization.
2.0.3. LLM Powered Reasoning for Generation
@@
The emergence of diffusion models has significantly advanced the field of image editing/manipulations (Wang et al., 2025). Recent progress in image inpainting, under both text-guided and unguided settings, has enabled more precise and high-quality manipulations (Wang et al., 2025). Text-conditioned inpainting (Xie et al., 2023) allows for detailed control over edits using natural language, while unconditioned methods offer strong baseline capabilities without explicit guidance. Traditionally, inpainting techniques (Huang et al., 2024) rely on the use of masks to specify regions for modification. In contrast, prompt-based image editing (Touvron et al., 2023a) allows for targeted alterations driven solely by text inputs, eliminating the need for explicit masks. Frameworks such as LANCE (Prabhu et al., 2023) and InstructPix2Pix (Brooks et al., 2022) harness this approach to automate image perturbation pipelines. LANCE (Prabhu et al., 2023), in particular, combines large language models (LLMs) with image captioning techniques to facilitate diverse, human-free image editing workflows. Expanding on this line of work, our method uses GPT-Image-1  (OpenAI, 2024), Gemini-2.0-Flash-Image-Generation (Anil and et al., 2023) and ICEdit (Zhang et al., 2025) to support image manipulation with a wider spectrum of perturbation intensities, guided by definitions of semantic change. To achieve this, we integrate models like Gemini-2.0-Flash (Anil and et al., 2023), ChatGPT-4o-latest (OpenAI, 2024) using prompt-based editing to generate precise, semantically meaningful image modifications. Multimodal models like Gemini, ChatGPT-4o, and Icedit leverage both text and image inputs to enable precise, context-aware edits and have a natural advantage over text-driven models like LANCE and InstructPix2Pix in fine-grained and semantically consistent image manipulation tasks. While text-driven models are effective for stylistic transformations, multimodal approaches offer broader capabilities and greater control over localized edits.
@@ 3.1. Data Generation Pipeline
(1) the person you identified appears naive
(2) the person you identified appears nonchalant
(3) the person you identified appears proud
(4) the person you identified appears remorseful
(5) the person you identified appears inexperienced
(6) some factual information depicted in the image changes
Do not include any other information in the response.
@@ 3.1.2. VLM based Image perception Manipulation
In this step, for each of the $<$referring expression, edit instruction$>$ pair, we prompt a VLM to perform the edit on the image, while ensuring that nothing else changes. We experiment with three different VLMs here - GPT-Image-1, Gemini-2.0-Flash-Image-Generation and ICEdit (Zhang et al., 2025). After observing  22K generated images, Gemini emerged as the VLM for highest quality manipulations, as its edits are the most coherent with the rest of the scene with no changes to the untargeted regions of the input image.The ICEdit model often results in easily identifiable fakes with significant artifacts, whereas GPT-image-1 tends to edit in a few cases, untargeted regions due to constraints in the output aspect ratios (current options are: $1024\times 1024$, $1536\times 1024$, $1024\times 1536$).
3.2. Analyzing the Manipulated images
@@
3.2.1. Ratio of edited area
To obtain a measure of the spatial degree of modification, we calculate the mean squared difference between the pixel values of the original and altered images, threshold the difference to remove noise, and perform connected component analysis to obtain the masks of the edited regions. The ratio of edited area to the total image area is then plotted over the entire dataset, shown in Figure 3(a), which displays a distribution spread out from 0 to 0.8, thus signifying the wide variety of edits present in our dataset.
3.2.2. Image Captioning
We use ShareGPT4V  (Chen et al., 2024) VLM model to obtain rich captions of both the original as well as tampered images. These are then used to calculate the directional similarity (Gal et al., 2022) between the real versus edited images and their corresponding captions as follows: ($CLIP_{OriginalImage}$ $CLIP_{TamperedImage}$) $\cdot$ ($CLIP_{OriginalImageCaption}$ $CLIP_{TamperedImageCaption}$). Particularly, we use Long-CLIP (Zhang et al., 2024) to calculate the image and text embeddings. From this analysis, we find that the maximum semantic change is observed in the object category, where the modification is around or on the person.
3.2.3. Level of manipulation performed
We categorize the manipulations into Person-level, Object-level and Scene-level depending upon whether the manipulation target is a person (facial expression, identity, gaze, pose, or clothing), an object connected to a person (i.e. an object in the foreground - appearance change, addition, removal etc.) or some component in the overall scene/background of the image respectively. Since an image can have multiple levels of changes depending upon the edit instruction, we display the distribution of the these alteration levels in our dataset through a Venn diagram in Figure 2(b). The distribution shows that the different levels of manipulations are widely distributed in our dataset with around one-third of the edits being person-only, around one-fifth being scene-only and one-sixth being object-only.
@@ 3.2.3. Level of manipulation performed
3.2.4. Perceptual Impact of the edit on the viewer
To obtain a measure of the change in the viewer perception as a result of the image tamperings, we consider the impact on the following six perceptual factors: (A) the viewer’s emotion or interpretation (B) the perceived identity or character of the person in the image (C) the perceived power dynamics/relationships (D) the perceived scene narrative (E) the possible intent of manipulation, and (F) the ethical implications of the tamperings. We prompt the Gemini-2.0-Flash VLM to describe these impacts, whose instruction details are shown in Prompt 3.2.3. The word clouds of the changes based on responses obtained from the VLM are shown in Figure 3. The first word map, corresponding to the changes in emotion perception (Figure 3(b)) features words such as significant, joyful, engaging, approachable, facial suggesting that our semantic modifications subtly influence the emotional perception. In case of scene context and narrative shifts (Fig. 3(c)), the word cloud contains terms such as focused, professional, potential, different, indicating plausible alterations to the conveyed story or setting due to the corresponding edits. Fig. 3(d), which illustrates the word map for shift in the perceived identity of depicted persons includes terms such asplayful, younger, vulnerable implying that our manipulations are able to influence the individual’s identity. The word map representing the intent behind the manipulations in Fig. 3(e) shows deceptive, persuasive and aesthetic as the most prominent words, reflecting a tendency for the manipulations to be viewed as intentional acts of persuasion or deception. According to the word map in Fig. 3(f), though the power dynamics and relationships remain stable for many cases, the presence of words like altered, dynamic, formal, competitive highlights that such dynamics do shift noticeably in a subset of the images. In Figure 4, we show some sample responses obtained for two such edits, with the different perceptual impacts highlighted. The VLM responses show that around 81% of the edits have mild ethical implications, while 14.2% have moderate and 0.3% of the edits have severe ethical implications in our dataset.
@@ 3.2.4. Perceptual Impact of the edit on the viewer
3.2.5. Quality of generations
To evaluate the visual quality of MultiFakeVerse, we used the peak signal-to-noise ratio (PSNR), structural similarity index (SSIM) (Wang et al., 2004) and Fréchet inception distance (FID) (Heusel et al., 2017) metrics as shown in Table 3. SSIM value 0 indicates no similarity between the compared images, while 1 indicates perfect similarity, therefore, our value of $0.5774$ is indicative of the targeted edits which ensure that the untargeted regions remain as close as possible to the original image. Similarly, a low FID value of 3.30 indicates that the quality and diversity of the generated fakes is excellent. A high PSNR value of 66.30 dB is indicative of good image quality.
@@ 3.2.7. Computational Cost
4. Benchmarks and Experiments
4.0.1. Data Partitioning and Evaluation Metrics
We split the dataset into train, validation, and test sets as follows: first we randomly select 70% of the real images as train set, 10% real images as validation set and remaining 20% as test set. Next we add the images generated corresponding to each of the real images in the same set as the real image. We evaluate detection using image-level accuracy and F1 scores. For forgery localization, we use Area Under the Curve (AUC), F1 scores and Intersection over Union (IoU).
4.0.2. Detection Evaluation
We compare MultiFakeVerse against several SOTA deepfake detection methods on the test set, including CnnSpot (Wang et al., 2020), AntifakePrompt (Chang et al., 2024), TruFor (Guillaro et al., 2023) and VLM based SIDA (Huang et al., 2024). To ensure a fair comparison, we first evaluate these models on our dataset using their original pre-trained weights (zero-shot setting), and then retrain two of them (CNNSpot and SIDA) with our train dataset to assess performance improvements. Table 4 demonstrates that these models trained on earlier inpainting-based fakes struggle to identify our VLM-Editing based forgeries, particularly, CNNSpot tends to classify almost all the images as real. AntiFakePrompt has the best zero-shot performance with 66.87% average class-wise accuracy and 55.55% F1 score. After finetuning on our train set, we observe a performance improvement in both CNNSpot and SIDA-13B, with CNNSpot surpassing SIDA-13B in terms of both average class-wise accuracy (by 1.92%) as well as F1-Score (by 1.97%).
4.0.3. Localization Evaluation
We evaluate the localization performance of SIDA-13B (Huang et al., 2024) model on our test set, using the perturbation masks obtained in Section 3.2, both using their pre-trained weights as well as after finetuning on our dataset. In zero-shot setting, the model has an IoU of 13.10, F1-Score 19.92 and AUC 14.06. After finetuning on our train dataset, the localization performance of the SIDA-13B model remains at an IoU of 24.74, F1-Score 39.40 and AUC 37.53, signifying that the model struggles to accurately identify the manipulated image regions. Thus, there is room for further improvement in the forgery localization methods, with MultiFakeVerse as an ideal benchmark.